OpenStudy (anonymous):

Can someone help me solve this expression: 15-20×-5+–1=

OpenStudy (anonymous):

Hey again :D one sec.

OpenStudy (ccswims):

remember PEMDAS

OpenStudy (anonymous):

Uh -9-20x, but if you are distributing then -9=20x = -2.222222222

OpenStudy (ccswims):

OH WAIT. lol i thought that "x" was a multiply symbol XD

OpenStudy (mathstudent55):

Is this the expression? \(\Large 15 - 20 \times (-5) + (–1) = \)

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

here is the math :P 15-20x-5+-1 15-5+-1=-9 -9-20x= +20x to both sides = -9=20x simple division leads us to -2.222222

OpenStudy (anonymous):

Wait no there aren't any parenthesis

OpenStudy (mathstudent55):

You must follow the correct order of operations. Multiplications and divisions come before additions and subtractions.

OpenStudy (anonymous):

simple isolation of the variable right ?

OpenStudy (anonymous):

Yeah @Etrainx

OpenStudy (calculusxy):

\[15 - 20 \times (-5) + (-1)\] \[15 + 100 + (-1)\] \[115 + (-1)\] \[114\]

OpenStudy (anonymous):

Should I multiply -20×-5 since its multiplication & following the order of pemdas ?

OpenStudy (anonymous):

OH i see

OpenStudy (anonymous):

you have 20 * (-5)

OpenStudy (anonymous):

i have 20x

OpenStudy (anonymous):

But the 20 is negative

OpenStudy (calculusxy):

The x really means multiplying. You're not working with variables, I believe.

OpenStudy (calculusxy):

@Halia You're correct. It's -20 because of the minus sign.

OpenStudy (anonymous):

when you subtract a negative number it becomes positive though.

OpenStudy (anonymous):

Yeah @calculusxy

OpenStudy (calculusxy):

@Etrainx You need to work with PEMDAS and this states that multiplying is before subtracting. Thus, you need to do -20 times -5.

OpenStudy (anonymous):

Wait so I multiped -20×-5 and I got +100

OpenStudy (anonymous):

I think these mathletes have it from here :P

OpenStudy (calculusxy):

Yes that's correct .

OpenStudy (mathstudent55):

That means you must start with \(20 \times (-5)\) \(\Large 15 - \color{red}{20 \times (-5)} + (–1) =\) \(\Large = 15 - \color{red}{(-100)} + (–1) \) Now we only have additions an subtractions, so we do them in the order they appear from left to right: \(\Large =15 - (-100) + (–1) \) \(\Large =15 + 100 -1 \) \(\Large =115 - 1\) \(\Large =114\)

OpenStudy (calculusxy):

The work that I did but explained in a better way @mathstudent55

OpenStudy (calculusxy):

;)

OpenStudy (calculusxy):

It's confusing, but just remember to take into consideration the sign (plus or minus) because that determines whether the following integer will be negative or positive.

OpenStudy (anonymous):

Mkay I understand it now :) thanks guys I really appreciate it !!!

OpenStudy (calculusxy):

You're very welcome @Halia :)

OpenStudy (mathstudent55):

Order of operations PEMDAS P - first do what is inside parentheses E - then do exponents MD - then do multiplications and divisions from left to right in the order they appear AS - finally do additions and subtractions in the order they appear from left to right

OpenStudy (mathstudent55):

You're welcome.